if roots of x 2 +bx+c=0 are real then prove that roots of x 2 +bx+c(x+a)(2x+b)=0 are also real.

if roots of x2+bx+c=0 are real then prove that roots of x2+bx+c(x+a)(2x+b)=0 are also real.

Grade:12

1 Answers

Vikas TU
7 years ago
x2+bx+c =  0
x
2+bx = -c ---------------------------(1)
x2+bx+c(x+a)(2x+b)=0 --------------------------(2)
put (1) in (2)
we get,
-c + c[2x^2 +(2a +b)x + ab -1] = 0 
c[2x^2 + (2a +b)x + ab -1] = 0
Now c cannot be equal to zero, 
therefore, 2x^2 + (2a +b)x + ab -1 = 0
and hence roots are real.